Three identical resistors of 4 ohms are connected to give a combined resistance of 6 ohms. Which of the diagrams illustrates how this was done?
1
Expert's answer
2020-04-20T10:01:48-0400

First two resistors will be in parallel. So,"\frac{1}{r}+\frac{1}{r}=\frac{2}{r}"

Now this resistor and the third resistor will be connected in series. So,

"\frac{r}{2}+r=\frac{3r}{2}"

Hence, the net resistance will be equal to "\frac{3r}{2}=6ohm"